OpenStudy (anonymous):

1.A drama club is planning a bus trip to New York City to see a Broadway play. The cost per person for the bus rental varies inversely as the number of people going on the trip. It will cost $30 per person if 44 people go on the trip. How much will it cost per person if 60 people go on the trip? Round your answer to the nearest cent, if necessary. A.$22.00 B.$40.91 C.$1,320.00 D.$21.29

OpenStudy (anonymous):

Let x = cost per person Let y = number of people >> The cost per person for the bus rental varies inversely as the number of people << y = k / x >> It will cost $30 per person if 44 people go on the trip << 44 = k / 30 k = 44 * 30 = 1320 y = 1320 / x >> How much will it cost per person if 60 people go on the trip? << 60 = 1320 / x x = 1320 / 60 = 22 <=====

OpenStudy (anonymous):

Help?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Yes thank you. sorry computer froze

OpenStudy (anonymous):

I miss you.